How many ways are there to assign 24 subjects into two equally sized groups?

The answer is not 5408312.
Answer #1

Number of ways to divide n students into k groups of size h is equal to

=\frac{n!}{h!^k *k!}

Here

n=24

h=12

k=2

=\frac{24!}{12!^2 *2!}=1352078